Lenows Drug Stores and Halls Pharmaceuticals are competitors in the discount drug chain store business. The separate capital structures for Lenow and Hall are presented here Lenow Hall Debt @ 9% $250,000 Debt @ 9% $500,000 Common stock, $10 par 500,000 Common stock, $10 par_250,000 Total $750,000 Total 750,000 25,000 Common shares 50,000 Common shares a. Complete the following table given earnings before interest and taxes of $29,000, $67,500, and $73,000. Assume the tax rate is 20 percent. (Negative amounts should be indicated by parentheses or a minus sign. Round your answers to 2 decimal places.) What is the relationship between Total assets Lenow- | EBIT/TAI%) EPS Hall EPS the EPS of th firms? $ 29,000 750,000 S67,500 $ 750,000 $ 73,000 750,000 b-1. What is the EBIT/TA rate when the firms have equal EPS? EBITITA rate b-2. What is the cost of debt? Cost of debt b-3. State the relationship between earnings per share and the level of EBIT. EPS is unaffected by financial leverage when the pre-tax return on assets (EBITITA the cost of debt. equals exceeds is less than c. If the cost of debt went up to 11 percent and all other factors remained equal, what would be the break-even level for EBIT? Break even level
Answer #1
a) What is the relationship between
EBIT [$] Total Assets [$] EBIT/TA [%] Lenow EPS Hall EPS the EPS of the two firms
29000 750000 3.87% 0.10 -0.51 Hall EPS < Lenow EPS
67500 750000 9.00% 0.72 0.72 Hall EPS = Lenow EPS
73000 750000 9.73% 0.81 0.90 Hall EPS > Lenow EPS
WORKINGS FOR N & EPS:
LENOW:
EBIT [$] Interest at 9% EBT Tax at 20% NI # of shares EPS
29000 22500 6500 1300 5200 50000 0.10
67500 22500 45000 9000 36000 50000 0.72
73000 22500 50500 10100 40400 50000 0.81
HALL:
EBIT [$] Interest at 9% EBT Tax at 20% NI # of shares EPS
29000 45000 -16000 -3200 -12800 25000 -0.51
67500 45000 22500 4500 18000 25000 0.72
73000 45000 28000 5600 22400 25000 0.90
b-1) EBIT/TA rate for equal EPS = 9%
b-2) Cost of debt = 9%
b-3) EPS is unaffected by financial leverage when the pre-tax return on assets (EBIT/TA) equals the cost of debt.
c) Break even level of EBIT is that EBIT level for which EPS = 0
EPS for Lenow = (E-27500)*0.80/50000
EPS for Hall = (E-55000)*0.80/25000
Where E = the break even EBIT.
Equating the two, we have
(E-27500)*0.80/50000 = (E-55000)*0.80/25000
Solving for E
(E-27500)/50 = (E-55000)/25
E-27500 = (E-55000)*2
E-27500 = 2*E-110000
110000-27500 = E = $82500
Break even level of EBIT = 82500
LENOW:
EBIT [$] Interest at 11% EBT Tax at 20% NI # of shares EPS
82500 27500 55000 11000 44000 50000 0.88
HALL:
EBIT [$] Interest at 11% EBT Tax at 20% NI # of shares EPS
82500 55000 27500 5500 22000 25000 0.88